A disk 254 mm in diameter accelerates uniformly from a speed of 1000 r/min to 2000

r/min in 20 s. Find (a) the angular acceleration in radians per second squared and (b) the

revolutions of the disk during the 20-s interval.


1
Expert's answer
2021-11-06T01:38:14-0400

(a)


"\alpha=\frac{\omega_2-\omega_1}{t}=0.83\text{ r/s}^2."

(b)


"N=\frac{\omega_2^2-\omega_1^2}{2\alpha}=502\text{ revolutions}"